John buys and sells phones. He bought phones at $950 each, and sold them (except for 5 that were fake and could not switch on) at $ 1120 each to make $850 profit.
Write the equation to solve this problem, and how many phones did John buy?

Answer by fractalier(6550)

Call the number of phones he bought, x.
Profit = Receipts - Costs
Plugging into that we get
850 = 1120(x-5) - 950x
Now solve this and we get
850 = 1120x - 5600 - 950x
6450 = 170x
x = 6450/170 = 37.94 or almost 38 phones